OpenStudy (anonymous):

A man was trying to swim to a buoy placed at a distance of 200 m out in a sea. it took him 1 min to swim 20 m . then a wave pushed him 10 m back and he rested for about 1 min brofre swimming again. he continued in this way for the rest of the journey . how long would it take the man to sim to the buo

OpenStudy (anonymous):

can u please tell me the solution

OpenStudy (anonymous):

tell me solution of this question

OpenStudy (anonymous):

My best guess after simple calculations would be 40 minutes. I'm pretty sure there would be some sort of way to calculate this but i used addition and subtraction.

OpenStudy (anonymous):

2 min for every 10m (2/x)=(10/200) proportion x=40